OpenStudy (anonymous):

Help me please!! At a playground, a 19.9-kg child plays on a slide that drops through a height of 2.31 m. The child starts at rest at the top of the slide. On the way down, the slide does a nonconservative work of -352 J on the child. What is the child's speed at the bottom of the slide?

OpenStudy (anonymous):

so we will set it up like this: PE_initial - 352J = KE_final make sense?

OpenStudy (anonymous):

Kind of. But where does the final speed come into play?

OpenStudy (anonymous):

KE = (1/2) m v^2, it comes in the KE_final, do you see the 'V' in the equation, it stands for velocity, aka, speed

OpenStudy (anonymous):

@jfisch1400 you still around?

OpenStudy (anonymous):

yeah. what about PE_initial?

OpenStudy (anonymous):

PE = mgh

OpenStudy (anonymous):

so this: PE_initial - 352J = KE_final now looks like this: mgh - 352 = (1/2)mv^2 you know mass, m, and g, gravity then just solve for v. i can help with the algebra if you need

OpenStudy (anonymous):

so, \[19.9*9.8*2.31-352=19.9v^2\] and i can do the algebra

OpenStudy (anonymous):

i got 2.225

OpenStudy (anonymous):

but that's not right.

OpenStudy (anonymous):

don't forget the (1/2) on the right hand side of the equation, like this: 19.9∗9.8∗2.31−352=(1/2) 19.9 v^2

OpenStudy (anonymous):

thank you very much. I have two others that I'm stuck on. could you help me with them?

OpenStudy (anonymous):

yep, just post them up and I will help you!

OpenStudy (anonymous):

I've been stuck on this one since last night. An 8.57-kg block slides with an initial speed of 1.70 m/s up a ramp inclined at an angle of 26.6° with the horizontal. The coefficient of kinetic friction between the block and the ramp is 0.64. Use energy conservation to find the distance the block slides before coming to rest.

OpenStudy (anonymous):

 so this one we will have to use sin(26.6) , but setting up the equation it will look like this: event 1 is the instant the block is sliding up at 1.7m/s event 2 is the moment the block comes to a stop PE + KE_1 -friction = KE + PE (this one is ugly!)

OpenStudy (anonymous):

PE_1 + KE_1 -friction = KE_2 + PE_2 mgh + (1/2)mv^2 - friction = (1/2)mv^2 + mgh

OpenStudy (anonymous):

and friction = μmgcosΘ right?

OpenStudy (anonymous):

but how do you get h_1 and h_2 for mgh?

OpenStudy (anonymous):

so we will say that PE_1 has h=0 becuase we are using this spot as our starting location. PE_1 = 0 KE_1 = (1/2)mv^2, we know it's mass, and we know it's veloocity so KE_1 = (1/2)8.57 * 1.7^2

OpenStudy (anonymous):

"and friction = μmgcosΘ right?" good, this is right, almost, inorder to use friction with KE and PE, we have to turn it into work, like this: Work = μmgcosΘ * distance

OpenStudy (anonymous):

"but how do you get h_1 and h_2 for mgh?" we set h_1 = 0, to use as our base elevation. h_2, is where we will need to use sin(26.6)

OpenStudy (anonymous):

this is tough stuff, what class is this? college?

OpenStudy (anonymous):

AP physics, yes. and okay

OpenStudy (anonymous):



OpenStudy (anonymous):

see how event 1 and event 2 represent the corners of the triangle?

OpenStudy (anonymous):

sin(26.6) = h_2/distance

OpenStudy (anonymous):

this part is crucial to solving this problem, is it making sense?

OpenStudy (anonymous):

now, this is where are are to now- KE_1 - work = PE_2 where KE_1 = (1/2)8.57 * 1.7^2 Work = μmgcosΘ * distance PE_2 = mgh_2

OpenStudy (anonymous):

brb, keep working on it, lemme know if you get stuck, but i think you are really close to getting this one!

OpenStudy (anonymous):

PE_1=0 KE_1=1/2(8.57)1.7^2=12.384 friction=.64(8.57)(9.8)d=53.751d KE_2=0 PE_2=(9.8(8.57))sin26.6=37.606 12.384-53.751d=37.606 -53.751d=25.222 d=.469 did I do evertything right? or did I miss anything?

OpenStudy (anonymous):

let me see.. on friction don't for get the cos() friction = .64*8.57*9.81*cos(26.6) * d and on PE_2, the d term PE_2 = 8.57*9.81*sin(26.6) * d

OpenStudy (anonymous):

so friction = 53.751cos(26.6)d = 48.062d and the equation is 12.384-48.062d = 37.606d 12.384=85.668d i got d = .145?

OpenStudy (anonymous):

I believe in you! haha

OpenStudy (anonymous):

yep, nicely done!

OpenStudy (anonymous):

d = 0.145 meters for me too

OpenStudy (anonymous):

you're the best. but I have one more if you don't mind?

OpenStudy (anonymous):

haha, one more!, throw it up, i'm good for a few more minutes before my gf calls

OpenStudy (anonymous):

An 81.0-kg in-line skater does +3550 J of nonconservative work by pushing against the ground with his skates. In addition, friction does -660 J of nonconservative work on the skater. The skater's initial and final speeds are 2.50 m/s and 1.60 m/s, respectively. (a) Has the skater gone uphill, downhill, or remained at the same level? Explain. (b) Calculate the change in height of the skater. (m)

OpenStudy (anonymous):

alright, so same thing as before, set up our equation: KE_1 + PE_1 + 3550j - 660j = KE_2 + PE_2

OpenStudy (anonymous):

in general we have: m = 81 g = 9.81 event 1 we have: v_1 = 2.5 event 2 we have: v_2 = 1.6 we need h_1 h_2 (if they exisit)

OpenStudy (anonymous):

ya?

OpenStudy (anonymous):

so what i think we should do is assume either h_1 = 0 or h_2 and when we solve for the term that's not 0, we can see if we assumed right. we can figure we assumed right if the sign on the h we are solving for comes out correct. if we picked the wrong h, then we can know it's wrong becuase when solving for the other h, the sign will be wrong - like lets say h should be -h, but when we solve it we get +h

OpenStudy (anonymous):

so if I get a positive h then it is uphill?

OpenStudy (anonymous):

yaya, IF you assume h_1 = 0, and if h_2 becomes positive, it means it's up hill

OpenStudy (anonymous):

sorry with the prior response, it was wordy!!

OpenStudy (anonymous):

@jfisch1400 I gotta take off soon, is there any more i can help?

OpenStudy (anonymous):

i got that h = .238 so he goes uphill? and this is the last problem i have. thank you for all the help!

OpenStudy (anonymous):

I got 3.82 with this equation .5*81*2.5^2 + 3550 -660 = .5*81*1.6^2 + 81 * h * 9.8

OpenStudy (anonymous):

assuming h_1 = 0, and solve for h_2

OpenStudy (anonymous):

oh I see what I did. Thank you so much!

OpenStudy (anonymous):

you're welcome. honestly, you are one of a few kids that actually learn, nice work!

OpenStudy (anonymous):

thanks. I'm always open to learn new things in the pursuit of knowledge and happiness. do you teach physics?

OpenStudy (anonymous):

I don't teach, but have been told that I should, huh.